OpenStudy (anonymous):

cosx-cos3x/sin3x-sinx = tan2x Show all your working please

OpenStudy (anonymous):

\[\frac{ cosx-\cos3x }{ \sin3x-sinx3 } = \tan2x\]

OpenStudy (anonymous):

Use two trigonometric identities for sin 3x and cos 3x. \[\sin(3x) = 3 \sin x - 4 \sin^3 x\] \[\cos(3x) = 4 \cos^3x - 3 \cos x\] Substitute in your formula \[\frac{ \sin x-\cos 3x }{ \sin 3x-\sin x } \] \[ = \frac{ \sin x - 4 \cos^3x + 3 \cos x } { 3 \sin x - 4 \sin^3 x - \sin x } \] \[ = \frac{ 4 \cos x (1- \cos^2x) } {2 \sin x (1-2 \sin^2x)} \] From trigonometric identities \[ 1= \sin^2 x + \cos^2x \] and \[ \cos 2x = 1-2 \sin^2x \] we get \[ = \frac{4 \cos x \sin^2x} {2 \sin x \cos 2x} \] From trigonometric identity \[ \sin 2x = 2 \sin x \cos x\] we get \[ = \frac {2 \sin x \sin 2x} {2 \sin x \cos 2x} = \tan 2x \]

OpenStudy (anonymous):

Thank you very much for the thorough explanation! :)

OpenStudy (anonymous):

@Steve.Irwin You are welcome:)